# Asynchronous-Many-Task Systems: Challenges and Opportunities - Scaling an AMR Astrophysics Code on Exascale machines using Kokkos and HPX

### v1

@@ 4.2 Kokkos
Kokkos implements a programming model in C++ for writing performance portable applications targeting all major HPC platforms. For that purpose, it provides abstractions for both parallel execution of code and data management. Kokkos is designed to target complex node architectures with N-level memory hierarchies and multiple types of execution resources. It currently can use CUDA, HIP, SYCL, HPX, OpenMP, and C++ threads as backend programming models [32]. With Kokkos, users can write their compute kernel once, then run it on the correct device by using the Kokkos execution space for this device and the associated Kokkos memory space for the data.
Furthermore, Kokkos also contains GPU-compatible SIMD types. On CPU these can be instantiated to use the appropriate SIMD instructions and registers of the current CPU, on GPU these can be instantiated to mere scalar values. This allows user to make explcit use of the SIMD ressources offered by the CPU, yet retain compatiblity with GPUs.
All these features further help to decouple the user code from specific hardware, greatly improving the portability.
@@ 4.3.1 Octo-Tiger’s Domain Science Use-Case:
Octo-Tiger is a distributed, full 3D, multi-scale, multi-model, adaptive mesh-refinement (AMR) astrophysics code based on FVM methods, designed to study stellar mergers as described in Section The Astrophysical Problem. It has been highly optimized for execution on distributed computing systems and can take advantage of the state of the art in GPUs. This allows Octo-Tiger to produce models with the high levels of resolution required to conduct convergence studies and understand the processes related to the increased $^{16}\textrm{O}$ concentration in DWD mergers that lead to the formation of R Coronae Borealis stars. Octo-Tiger is a real-world application that has proven to solve pressing science domain problems in the field of simulating the merger of two white dwarf stars bound in a binary star system [33, 5, 17].
The Octo-Tiger code has several features that make it particular suited for modelling interacting binaries. The entire grid structure rotates with the initial orbital frequency of the binary. This reduces the velocity of the material relative to the grid, and therefore reduces the non-physical effects due to numerical viscosity. The code carefully conserves global quantities like linear momentum, angular momentum, and energy, which is important for modelling the initial phase of mass transfer as small violations in conservation of those quantities can significantly effect the simulation results. Octo-Tiger conserves these quantities in the hydrodynamics solver by matching fluxes across jumps in refinement levels. As is the case with most FMM solvers, the FMM conserves linear momentum. We also developed a special technique that enables the FMM to simultaneously conserve angular momentum. While the hydrodynamics solver does not conserve angular momentum, conservation in the gravity solver enables Octo-Tiger to conserve energy in the rotating frame; otherwise this would only be possible in a non-rotating frame. The AMR grid allows for the extension of the spatial domain boundaries to orders of magnitude larger than the orbital separation, allowing Octo-Tiger to model the extended low density outflows that occur during the merger process. The AMR grid also enables higher resolution in the regions of interest, making convergence studies more feasible.
@@ 4.3.2 Octo-Tiger’s Components and Data-Structure:
Octo-Tiger models binary star systems as self-gravitating fluids. Hence, it uses two interleaved solvers: It employs a Fast Multipole Method (FMM) gravity solver and a finite volume hydrodynamics solver. A third, experimental solver for radiation is currently in active development. As Octo-Tiger uses a third-order runge kutta integration scheme, each time step involves three iterations of the hydrodynamics solver. As the FMM is modified to conserve angular momentum, each time step further involves actually six (instead of three) iterations of the FMM solver.
The solvers operate on an adaptive octree, with the AMR focusing on the atmosphere between the stars where the mass exchange between them is happening. Each tree node contains an entire sub-grid with ${8}\times{8}\times{8}$ ($512$) cells to improve the computational efficiency. All compute kernels usually operate on one such sub-grid (and its ghost layers) at a time, meaning we deal with lots of small, potentially concurrent compute kernels, but each of them only contains a comparatively small compute load. While we can adjust the size of the sub-grids at compile-time to compensate, this would negatively impact the adaptivity (as we would get a less refined grid given the same amount of overall cells), scalability (as the sub-grids are the components we distribute onto the compute-nodes) and lastly, it would impact the runtime of the FMM (as it uses the tree-structure to avoid computations by approximations).
@@ 5 Resolving the Challenges of Scalable AMR Applications with HPX and Kokkos
Using HPX in concert with Kokkos presents the oppurtunity to easily develop portable, yet highly scalable, distributed codes, even when using irregular tree-based structures. Thus, we turned to both frameworks when developing Octo-Tiger and porting it to GPUs, especially given our small team of core developers with only one developer being available for the refactoring and porting effort.
@@
  • We needed to improve the HPX-Kokkos interoperability, avoiding both, the blocking of HPX worker threads with Kokkos fences and avoiding conflicting thread pools for host execution.
  • We needed to address the issue of GPU device starvation caused by compute kernels that are too small (in turn caused by the small workloads per tree-node in AMR applications such as Octo-Tiger). We furthermore needed to resolve the memory allocation overheads caused by temporary (but necessary) GPU buffers.
@@
In this section, we outline how we addressed each of those concerns in our previous work. Thus, we introduce the HPX-Kokkos integrations (intended to address the HPX-Kokkos interoperability) [1], the CPPuddle allocators and executors (intended to address work starvation and memory overheads for temporary buffers) [2], the SVE and RVV SIMD types (to enable portable vectorization for A64Fx and RISC-V CPUs) [3] and the LCI parcelport (offering another networking choice within HPX) [6].
@@ 5.1 Combining Kokkos and HPX
The first integration results in HPX-Kokkos that enables the execution of asynchronous Kokkos Kernels (or deep copies) into the HPX task graph. This makes it easy to define continuation tasks that should be triggered once a kernel is done executing, such as doing post-processing on the results or communicating them, easing the development of distributed code. This also eliminates any need for a CPU thread to actively wait on some GPU kernel, as long as there is some other work remaining in the queue, allowing the runtime to easily manage thousands of concurrent kernel launches, using hundreds of GPU streams and just a few CPU worker threads on each process (usually one thread per utilized core).
HPX-Kokkos itself is a thin compatibility layer: depending on which Kokkos execution space is used, it uses the appropriate HPX API integrations to get an HPX Future for this execution space that represents the event of the encapsulated work being done executing. For instance, when using a Kokkos CUDA execution space, HPX-Kokkos will use the get_future functionality of the underlying HPX-CUDA integration that is directly implemented within HPX using event polling. Respectively, it will use the appropriate integrations for the Kokkos HIP and Kokkos SYCL execution spaces as well (as HPX includes HIP and SYCL integrations which we developed previously [1, 34]). HPX-Kokkos merely maps the calls correctly and provides convenience functions such as deep_copy_async or parallel_for_async that call their Kokkos equivalents (deep_copy or parallel_for) and immediately obtain a HPX Future for those.
@@
Using Kokkos and HPX together with these integrations allows us to both take full advantage of HPX (allowing us to finely interleave computation and communication tasks while transparently managing asynchronous execution graphs) and Kokkos (allowing us to target different CPU and GPU hardware with a single kernel implementation).
@@ 5.2 GPU-compatible SIMD Implementation with SVE Support
While using Kokkos itself already allows us to target both CPU and GPU machines with the same kernel implementation, we still need to ensure that the kernel makes use of SIMD vectorization on the CPU. Although relying on autovectorization might work in some kernels, the ones in Octo-Tiger contain too many branches, preventing the compiler from vectorizing the code. We can work around this problem by using SIMD masking. However, we still want to do this in a portable way as the code still needs to execute correctly on the GPU.
For this reason, we use C++ SIMD datatypes, specifically we use a common subset of the Kokkos SIMD types and the std::simd types (allowing us to decide at compile time, which of those to apply). These Kokkos SIMD types already overload all relevant operators and math operations, allowing us to use them just like normal floating point types. However, during compile time they get either instantiated based on the relevant SIMD intrinsics (such as AVX512 or SVE) or as scalar floating point types for kernels that run on the GPU. These types also support SIMD masks for conditional code branching.
We added the SIMD types and masks in all major Octo-Tiger compute kernels and tested the SIMD speedup in [3]. We further tested instantiating the types with std::simd types and introduced our own implementation of those for SVE in that work, in order to prepare for our Fugaku runs.
Together with the Kokkos HPX execution space, this allows us to have efficient CPU kernel implementations. Usually, we use rather fine-grained kernels where we utilize just 1 to 16 worker threads per kernel, instead of the entire CPU. With many of those kernels running concurrently, it allows us to finely interleave computation and communication, while at the same time utilizes the SIMD capabilities of the CPU to the best extent possible. We implemented SIMD types similar to those for RISC-V as well [4].
@@ 5.3 Addressing GPU Starvation and Memory Overheads
While fine-grained compute kernels (as the ones within Octo-Tiger) can be beneficial for adaptivity (more tree levels) and scalability (finer tree structure to distribute onto the compute nodes), it presents major problems for a GPU implementation: On a GPU, we not only have more compute elements than on an average CPU, but also rely on having enough work items to hide latencies and stalls. At worst, a too small GPU compute kernel with few work items will cause outright device starvation (where we do not even scale to all SMs on an NVIDIA A100). In addition, even when there is enough work for all streaming multiprocessors/compute units, the kernel will still run inefficiently if there are not enough concurrent work items available for latency hiding.
@@
Handling concurrent GPU kernel launches is straightforward thanks to the integration of Kokkos kernel execution into the HPX task graph via HPX-Kokkos. Hence, we can easily have $128$ concurrent kernels per process, even if the said process only contains a few worker threads to handle launches and communications. However, the overhead of creating a GPU stream (and thus a GPU executor) is significant, and hence we use a pre-allocated pool of GPU executors in each process, combined with a scheduler drawing from it on demand.
To enable the dynamic kernel fusion within HPX we introduced a special executor in [2]. These executors (enabled by HPX Futures) allow users to define code regions where multiple HPX tasks can cooperate by concurrently fusing their individual GPU kernels into one larger kernel (provided it is the same compute kernel operating on different data items). This reduces the overall number of GPU related API calls (fewer kernels and data transfers need to be initiated) and enlarges the work size of the fused kernel (better utilizing the GPU). This is especially valuable on GPUs that are less capable of efficiently running concurrent individual kernels.
We compared these three techniques to avoid GPU starvation and their impact within the Octo-Tiger hydro solver in greater detail in [2]. Here we also compared the performance of the native CUDA/HIP kernels and their Kokkos counterparts on both NVIDIA A100 and AMD MI100 GPUs.
Lastly, when dealing with many fine-grained compute kernels and associated communication ghost layers, the amount of required temporary GPU buffers can become a problem. Allocating and de-allocating them on-the-fly is too expensive, but pre-allocating them for each tree node would significantly increase the memory requirements of Octo-Tiger. We avoid this problem by developing a dynamically growing memory pool for GPU device memory and for pinned memory on the host side. If we need a temporary GPU buffer, we can allocate it with a special allocator that will recycle a (currently unused) GPU buffer from the memory pool. Upon de-allocation, the buffer will simply go back into the pool of unused memory (but stays allocated from the perspective of the system). If there is no unused buffer of sufficient size available upon allocation, a new one will be created. This way, we only allocate as much memory for the temporary and communication buffers as is needed for maximum concurrency, allowing us to run larger scenarios with fewer compute nodes.
@@ 5.4 Inter-process Communication Optimizations
Adding on top of the techniques used in the MPI parcelport, the LCI parcelport [6] further employs many communication features that do not exist in current standard MPI, including (a) better communication primitives for active message style communication to eliminate tag matching, ordering, and memory copy overhead; (b) lock-free completion queues for efficient polling of a large number of pending communication operations; (c) lightweight interaction with and replication of low-level communication resources to expose more parallelism from the shardware level. Together, these techniques enhance the HPX communication layer with the ability to more efficiently handle a larger number of (possibly small) concurrent communications in a multi-threaded environment.

### v2

@@ 4.2 Kokkos
Kokkos implements a programming model in C++ for writing performance portable applications targeting all major HPC platforms. For that purpose, it provides abstractions for both parallel execution of code and data management. Kokkos is designed to target complex node architectures with N-level memory hierarchies and multiple types of execution resources. It currently can use CUDA, HIP, SYCL, HPX, OpenMP, and C++ threads as backend programming models [40]. With Kokkos, users can write their compute kernel once then run it on the correct device by using the Kokkos execution space for the device and the associated Kokkos memory space for data.
Furthermore, Kokkos also contains GPU-compatible SIMD types. On CPU, these can be instantiated to use the appropriate SIMD instructions and registers of the current CPU. On GPU, these can be instantiated to mere scalar values. This allows user to make explicit use of the SIMD resources offered by the CPU, yet retain compatibility with GPUs.
These features help to decouple user code from specific hardware, greatly improving the portability.
@@ 4.3.1 Octo-Tiger’s Domain Science Use-Case:
Octo-Tiger is a distributed, full 3D, multi-scale, multi-model, adaptive mesh-refinement (AMR) astrophysics code based on FVM methods and designed to study stellar mergers as described in Section The Astrophysical Problem. It has been highly optimized for execution on distributed computing systems and can take advantage of the state of the art in GPUs. This allows Octo-Tiger to produce models with the high levels of resolution required to conduct convergence studies and understand the processes related to the increased $^{16}\textrm{O}$ concentration in DWD mergers that lead to the formation of R Coronae Borealis stars. Octo-Tiger is a real-world application that has proven to solve pressing science domain problems in the field of simulating the merger of two white dwarf stars bound in a binary star system [41, 5, 25].
The Octo-Tiger code has several features that make it particularly well-suited for modeling interacting binaries. The entire grid structure rotates with the initial orbital frequency of the binary. This reduces the velocity of the material relative to the grid, and therefore reduces the non-physical effects due to numerical viscosity. The code carefully conserves global quantities like linear momentum, angular momentum, and energy, which is important for modeling the initial phase of mass transfer as small violations in conservation of those quantities can significantly effect the simulation results. Octo-Tiger conserves these quantities in the hydrodynamics solver by matching fluxes across jumps in refinement levels. As is the case with most Fast Multipole Method (FMM) solvers, the FMM conserves linear momentum. We also developed a special technique that enables the FMM to simultaneously conserve angular momentum. While the hydrodynamics solver does not conserve angular momentum, conservation in the gravity solver enables Octo-Tiger to conserve energy in the rotating frame; otherwise this would only be possible in a non-rotating frame. The AMR grid allows for the extension of the spatial domain boundaries to orders of magnitude larger than the orbital separation. This allows Octo-Tiger to model the extended low density outflows that occur during the merger process. The AMR grid also enables higher resolution in the regions of interest, making convergence studies more feasible.
@@ 4.3.2 Octo-Tiger’s Components and Data-Structure:
Octo-Tiger models binary star systems as self-gravitating fluids. Hence, it uses two interleaved solvers: It employs a Fast Multipole Method (FMM) gravity solver and a finite volume hydrodynamics solver. A third, experimental solver for radiation is currently in active development. As Octo-Tiger uses a third-order Runge-Kutta integration scheme, each time step involves three iterations of the hydrodynamics solver. As the FMM is modified to conserve angular momentum, each time step further involves actually six (instead of three) iterations of the FMM solver.
The solvers operate on an adaptive octree with the AMR focusing on the atmosphere between the stars, which is where the mass is being exchanged between stars. Each tree node contains an entire sub-grid with ${8}\times{8}\times{8}$ ($512$) cells to improve the computational efficiency. Each compute kernel typically operates on one such sub-grid (and its ghost layers) at a time. As a result, we deal with a large number of potentially concurrent compute kernels when traversing the tree. However, each of these kernels only processes a comparatively small compute load. In theory, we can further increase the size of the sub-grids at compile-time to compensate, thus increasing the compute load per individual kernel invocation for better efficiency. Yet, this would in turn negatively impact the adaptivity (as we would get a less refined grid given the same amount of overall cells), the scalability (as the sub-grids are the components we distribute onto the compute-nodes) and lastly, it would impact the runtime of the FMM (as it uses the tree-structure to avoid computations by approximations).
@@ 5 Resolving the Challenges of Scalable AMR Applications with HPX and Kokkos
Using HPX in conjunction with Kokkos presents the opportunity to easily develop portable, yet highly scalable, distributed codes, even when using irregular tree-based structures. For this reason, we turned to both frameworks when developing Octo-Tiger and porting it to GPUs, especially given our small team of core developers with only one developer being available for the refactoring and porting effort.
@@
  • We needed to improve the HPX-Kokkos interoperability, avoiding both the blocking of HPX worker threads with Kokkos fences and avoiding conflicting thread pools for host execution.
  • We needed to address the issue of GPU device starvation caused by compute kernels that are too small (in turn, caused by the small workloads per tree-node in AMR applications such as Octo-Tiger). We furthermore needed to resolve the memory allocation overheads caused by temporary (but necessary) GPU buffers.
@@
In this section, we outline how we addressed each of those concerns in our previous work. Thus, we introduce the HPX-Kokkos integrations (intended to address the HPX-Kokkos interoperability) [1], specialized allocators and executors (intended to address work starvation and memory overheads for temporary buffers) [2], additional SIMD types (to enable portable vectorization for A64FX and RISC-V CPUs while maintaining GPU compatibility of the kernels) [3] and the LCI parcelport (offering another networking choice within HPX) [6].
@@ 5.1 Combining Kokkos and HPX
The first integration resulted in the HPX-Kokkos interoperability library, which enables the execution of asynchronous Kokkos Kernels (or deep copies) into the HPX task graph. This makes it easy to define continuation tasks that should be triggered once a kernel is done executing, such as doing post-processing on the results or communicating them, easing the development of distributed code. This also eliminates the need for a CPU thread to actively wait on some GPU kernel as long as there is some other work remaining in the queue. This allows the runtime to easily manage thousands of concurrent kernel launches, using hundreds of GPU streams and just a few CPU worker threads on each process (usually one thread per utilized core).
HPX-Kokkos itself is a thin compatibility layer: depending on which Kokkos execution space is used, it uses the appropriate HPX API integrations to get an HPX Future for this execution space that represents the event of the encapsulated work being done executing. For instance, when using a Kokkos CUDA execution space, HPX-Kokkos will use the get_future functionality of the underlying HPX-CUDA integration that is directly implemented within HPX using event polling. Similarly, it will use the appropriate integrations for the Kokkos HIP and Kokkos SYCL execution spaces as well (as HPX includes HIP and SYCL integrations which we developed previously [1, 42]). HPX-Kokkos merely maps the calls correctly and provides convenience functions such as deep_copy_async or parallel_for_async that call their Kokkos equivalents (deep_copy or parallel_for) and immediately obtain a HPX Future for those.
@@
Using Kokkos and HPX together with these integrations allows us to both take full advantage of HPX (allowing us to finely interleave computation and communication tasks while transparently managing asynchronous execution graphs) and Kokkos (allowing us to target different CPU and GPU hardware with a single kernel implementation). Figure 3 shows an example of how a part of such a graph would look like for Octo-Tiger.
Figure 3: Illustration of a small part of Octo-Tiger’s task-graph, depicting the tasks that are appended to the task-graph for every sub-grid in every iteration of Octo-Tiger’s hydro solver. Tasks (in white) are represented with HPX futures (for instance, named com_fut). These futures can be dynamically and asynchronously chained together into a task-graph using HPX functionality (in gray), such as then and when_all. This allows us to build the task-graph on-the-fly whilst traversing Octo-Tiger’s tree. With the HPX-Kokkos integration, it is possible to integrate asynchronous Kokkos kernels (in green) into this task-graph as well, freeing us from having to use blocking fences to process their results. Overall, this results in a multitude of concurrent CPU tasks, GPU tasks, and communication tasks. This leads to an interleaving of tasks, helping us to better utilize the available hardware resources and achieve scalability.
@@ 5.2 GPU-compatible SIMD Implementation with SVE Support
While using Kokkos itself already allows us to target both CPU and GPU machines with the same kernel implementation, we still need to ensure that the kernel makes use of SIMD vectorization on the CPU. Although relying on auto-vectorization might work in some kernels, the ones in Octo-Tiger contain too many branches, preventing the compiler from automatically vectorizing the code. We can work around this problem by using SIMD masking. However, we still want to do this in a portable way as the code still needs to execute correctly on the GPU.
For this reason, we use C++ SIMD datatypes, specifically we use a common subset of the Kokkos SIMD types and the std::simd types (allowing us to decide at compile time, which of those to apply). As our work predates C++26 (and the official introduction of std::simd) we used the std::experimental::simd implementation within GCC (version 11 and upwards) during our development and testing. However, this is compatible with std::simd. Both the Kokkos SIMD types and the std::experimental::simd overload all relevant operators and math operations, allowing us to use them just like normal floating point types. During compile time, they get either instantiated based on the relevant SIMD intrinsics (such as AVX512) or as scalar floating point types for kernels that run on the GPU. The types also support SIMD masks for conditional code branching.
We added the SIMD types and masks in all major Octo-Tiger compute kernels and tested the SIMD speedup in [3] both with the Kokkos SIMD types and with std::experimental::simd types from GCC. Furthermore, as there were no std::simd-compatible types available for SVE, we implemented these types ourselves and tested them within Octo-Tiger. This allows us to also target the SIMD units in ARM processors (particularly, the A64FX CPUs used in the Supercomputer Fugaku). Eventually, our SVE SIMD type implementation from  [3] has been adapted for inclusion into GCC to benefit the wider HPC community  [^2] Merged into GCC in https://github.com/gcc-mirror/gcc/commit/9ac3119fec81fb64d11dee8f853145f937389366. Similar to the implementation of the SVE types, we also implemented SIMD types for RISC-V and its RVV SIMD instructions [4]. Unlike our SVE types, these have not been yet merged into GCC yet.
Together with the Kokkos HPX execution space, this allows us to have efficient CPU kernel implementations. Usually, we use rather fine-grained kernels where we utilize just 1 to 16 worker threads per kernel instead of the entire CPU. With many of those kernels running concurrently, it allows us to finely interleave computation and communication while at the same time utilizing the SIMD capabilities of the CPU to the best extent possible.
@@ 5.3 Addressing GPU Starvation and Memory Overheads
While fine-grained compute kernels (as the ones within Octo-Tiger) can be beneficial for adaptivity (more tree levels) and scalability (finer tree structure to distribute onto the compute nodes), it presents major problems for a GPU implementation: On a GPU, we not only have more compute elements than on an average CPU, but we also rely on having enough work items to hide latencies and stalls. At worst, a too small GPU compute kernel with few work items will cause outright device starvation (where we do not even scale to all SMs on an NVIDIA A100). In addition, even when there is enough work for all streaming multiprocessors/compute units, the kernel will still run inefficiently if there are not enough concurrent work items available for latency hiding.
@@
Handling concurrent GPU kernel launches is straightforward thanks to the integration of Kokkos kernel execution into the HPX task graph via HPX-Kokkos. With this integration, we can easily have $128$ concurrent kernels per process, even if the said process only contains a few worker threads to handle launches and communications. However, the overhead of creating a GPU stream (and thus a GPU executor) is significant. To avoid this overhead, we use a pre-allocated pool of GPU executors in each process, combined with a scheduler drawing from it on demand.
To enable the dynamic kernel fusion within HPX, we introduced a special executor in [2]. These executors (enabled by HPX Futures) allow users to define code regions where multiple HPX tasks can cooperate by concurrently fusing their individual GPU kernels into one larger kernel (provided it is the same compute kernel operating on different data items). This reduces the overall number of GPU related API calls (fewer kernels and data transfers need to be initiated) and enlarges the work size of the fused kernel (better utilizing the GPU). This is especially valuable on GPUs that are less capable of efficiently running concurrent individual kernels.
We compared these three techniques to avoid GPU starvation and their impact within the Octo-Tiger hydro solver in greater detail in [2]. Here, we also compared the performance of the native CUDA/HIP kernels and their Kokkos counterparts on both NVIDIA A100 and AMD MI100 GPUs.
Lastly, when dealing with many fine-grained compute kernels and associated communication ghost layers, the amount of required temporary GPU buffers can become a problem. Allocating and de-allocating them on-the-fly is too expensive but pre-allocating them for each tree node would significantly increase the memory requirements of Octo-Tiger. We avoid this problem by developing a dynamically growing memory pool for GPU device memory and for pinned memory on the host side. If we need a temporary GPU buffer, we can allocate it with a special allocator that will recycle a (currently unused) GPU buffer from the memory pool. Upon de-allocation, the buffer will simply go back into the pool of unused memory (but stays allocated from the perspective of the system). If there is no unused buffer of sufficient size available upon allocation, a new one will be created. This way, we only allocate as much memory for the temporary and communication buffers as is needed for maximum concurrency, allowing us to run larger scenarios with fewer compute nodes.
Both the kernel fusion functionality and the memory pool functionality are exposed through specialized executors and allocators which we published in the utility library CPPuddle.
@@ 5.4 Inter-process Communication Optimizations
Adding on top of the techniques used in the MPI parcelport, the LCI parcelport [6] further employs many communication features that do not exist in current standard MPI, including (a) better communication primitives for active message style communication to eliminate tag matching, ordering, and memory copy overhead; (b) lock-free completion queues for efficient polling of a large number of pending communication operations; (c) lightweight interaction with and replication of low-level communication resources to expose more parallelism from the hardware level. Together, these techniques enhance the HPX communication layer with the ability to more efficiently handle a larger number of (possibly small) concurrent communications in a multi-threaded environment.